# Gotta Hear Them All: Sound Source-Aware Vision to Audio Generation

### v2

@@ A.2.1 Cross-Modal Sound Source Manifold
We filter source-unannotated visual-audio pairs from VGGSound [3] with an open-vocabulary object segmentor, CLIP as RNN (CaR) [52], keeping the pairs where only one visual region is segmented. The confidence threshold of CaR is set to 0.5. We use the VGGSound category labels as segmentation vocabulary. CaR’s pixel-level segmentations are abstracted into bounding boxes to capture fuller visual content. We crop each video’s central frame with the predicted bounding box to pair with its audio clip and verify the data quality by manually reviewing 10 results from each category.
@@ A.2.2 Sound Source Remixer
Recall Eq. 8. We assign a learned [cls] token at the head of each token sequence for the Sound Source Remixer’s prediction. The tokens first travel through a stack of attention modules, where each module contains an Efficient Attention layer followed by a feed forward network and an ELU activation. The [cls] token is then passed to two MLP heads to respectively estimate the mean and variance of the mixed CLAP audio embeddings. We then sample these embeddings from a normal distribution with these estimated parameters. The embeddings are further normalized with respect to their $l2$-norms to respect the original representation format of CLAP. Each MLP head is three-layer with [768, 640, 512] neurons and ELU activations. We use only one Efficient Attention layer following the optimal setup from ablation experiments in Sec. A.4.3.
@@ A.3.2 Modality Alignment Tests
We randomly select 4 audio categories and conduct a discriminant test on each in Tab. 6, which is also visualized as a canonical plot in Fig. 8. The discriminant model is a wide linear binary classifier to predict whether a given sample is from the visual or audio modality. We observe that this classifier works perfectly on raw embedding samples but fails to classify the CMSS embeddings with a low Entropy $R^{2}$ (classification contingency) and high -2 log-likelihood (classification uncertainty). The discriminant test supports the manifold’s ability to close the modality gap between visual and audio data, as the discriminant classifier is significantly confused after the manifold remapping.
@@ A.4.1 Ablation of Cycle Mix
We record the ablations of Cycle Mix in Algorithm 1 with Tab. 8. When the iteration is 1 and sampling size is 1, we directly obtain the Remixer’s output without Cycle Mix. Higher sample size and iterations lead to better multi-source generations while single-source performance slightly drops. Setting the parameters too high compromises performance in both generation tasks. Since this work’s primary focus is multi-sound-source V2A synthesis, we determine a sample size of 64 and an iteration count of 64 as the optimal parameter setup and use it throughout other experiments.
@@ Appendix A.5 Setup of Subjective Evaluation
We disseminate an online survey for the subjective evaluation and collect results from 20 participants to measure generation fidelity and relevance of our method along with baselines. The baseline methods include Im2Wav, Diff-Foley-Image, S&H-Text, and V2A-Mapper as described in Appendix A.1. In the first survey section, we ask the participants to sign a consent form as illustrated in Fig. 9 (a). The non-consenting participants are screened out without any data collection. In the second section, we ask 20 fidelity-rating questions without visual cues following Fig. 9 (b). To unify the comparison context, we include a short tag in the question describing the ground-truth audio content. For each generated sample, the testee is asked to give out fidelity rating on a 1-5 scale. In the third section, we ask 20 relevance-rating questions given the visual condition used during generation, which is deicted in Fig. 9 (c). The ratings are also collected on a 1-5 scale.
@@ Appendix A.6 Implementation
We attach the implementation code of SSV2A with this supplementary material, which can be found at https://github.com/wguo86/SSV2A. The ssv2a python module includes both training and evaluation implementations necessary for our SSV2A tasks. Please see the readme.md file in root directory for specific execution instructions.

### v3

@@ A.2.1 Cross-Modal Sound Source Manifold
We filter source-unannotated visual-audio pairs from VGGSound [3] with an open-vocabulary object segmentor, CLIP as RNN (CaR) [55], keeping the pairs where only one visual region is segmented. The confidence threshold of CaR is set to 0.5. We use the VGGSound category labels as segmentation vocabulary. CaR’s pixel-level segmentations are abstracted into bounding boxes to capture fuller visual content. We crop each video’s central frame with the predicted bounding box to pair with its audio clip and verify the data quality by manually reviewing 10 results from each category. The resulted VGGS3 dataset has 106514 samples across 221 sound source categories, which promises audio diversity. One potential bias is that the curated sound sources have unbalanced category frequencies with a max of 946 and a min of 100. We intend to release a balanced version of VGGS3 alongside the current version.
@@ A.2.2 Sound Source Remixer
Recall Eq. (8). We assign a learned [cls] token at the head of each token sequence for the Sound Source Remixer’s prediction. The tokens first travel through a stack of attention modules, where each module contains an Efficient Attention layer followed by a feed forward network and an ELU activation. The [cls] token is then passed to two MLP heads to respectively estimate the mean and variance of the mixed CLAP audio embeddings. We then sample these embeddings from a normal distribution with these estimated parameters. The embeddings are further normalized with respect to their $l2$-norms to respect the original representation format of CLAP. Each MLP head is three-layer with [768, 640, 512] neurons and ELU activations. We use only one Efficient Attention layer following the optimal setup from ablation experiments in Sec. A.4.3.
@@ A.3.2 Modality Alignment Tests
We randomly select 4 audio categories and conduct a discriminant test on each in Tab. 7. These results are also visualized as a canonical plot in Fig. 8. The discriminant model is a wide linear binary classifier to predict whether a given sample is from the visual or audio modality. We observe that this classifier works perfectly on raw embedding samples but fails to classify the CMSS embeddings with a low Entropy $R^{2}$ (classification contingency) and high -2 log-likelihood (classification uncertainty). This discriminant test supports the manifold’s ability to close the modality gap between visual and audio data distributions, as the discriminant classifier is significantly confused after the manifold remapping of CLIP and CLAP embeddings.
@@ A.4.1 Ablation of Cycle Mix
The Cycle Mix algorithm has two adjustable parameters: the number of iterations and the the sampling size of remixed CLAP embeddings in each iteration. We conduct 8 VGG-SS tests to observe the effect of Cycle Mix parameters. These ablations are illustrated in Tab. 9. When the iteration is 1 and sampling size is 1, we directly obtain the Remixer’s output without Cycle Mix. Higher sample size and iterations lead to better multi-source generations while single-source performance slightly drops. Setting the parameters too high compromises performance in both generation tasks. Since this work’s primary focus is multi-sound-source V2A synthesis, we determine a sample size of 64 and an iteration count of 64 as the optimal parameter setup and use it throughout other experiments.
@@ Appendix A.5 Setup of Subjective Evaluation
We disseminate an online survey for the subjective evaluation and collect results from 20 participants to measure generation fidelity and relevance of our method along with baselines. The baseline methods include Im2Wav, Diff-Foley-Image, S&H-Text, and V2A-Mapper as described in Appendix A.1. In the first survey section, we ask the participants to sign a consent form as illustrated in Fig. 12 (a). The non-consenting participants are screened out without any data collection. In the second section, we ask 20 fidelity-rating questions without visual cues following Fig. 12 (b). To unify the comparison context, we include a short tag in the question describing the ground-truth audio content. For each generated sample, the testee is asked to give out fidelity rating on a 1-5 scale. In the third section, we ask 20 relevance-rating questions given the visual condition used during generation, which is depicted in Fig. 12 (c). The ratings are also collected on a 1-5 scale.
Appendix A.6 Why Not Cascaded Composition?
The most straightforward way to composite multimodal sound sources into a single audio is to generate an audio track for each source condition via video-to-audio or text-to-audio models and overlay them together. However, such a cascading audio synthesis system lacks interaction, context and style awareness when integrating multiple sound sources, which are keys to a convincing audio scene. We show with qualitative examples in this section that our SSV2A composition achieves these features. These examples are best experienced on our website at the Composition Comparisons section.
A.6.1 Interaction Awareness
Figure 9: Example of interaction awareness. Our composition arranges drum and bass sounds into interactive music.
We generate a drum-only audio clip and a bass-only clip with SSV2A. Simply overlaying these clips yields a mixed track with no interactions between these instruments, which is shown in Fig. 9. With our SSV2A composition, we are able to generate an appealing piece of drum-bass music with rich interactions.
A.6.2 Context Awareness
Figure 10: Example of context awareness. Our composition transforms a normal talking man into a police officer on duty.
We generate audio clips of a normal speech and an academic conference in Fig. 10. Cascaded composition synthesizes an audio clip with conflicting talkers and no conference room reverb. Our SSV2A composition properly transfers the speech style into a reverberated academic presentation.
A.6.3 Style Awareness
Figure 11: Example of style awareness. Our composition changes a normal speech into an academic presentation with conference room reverb.
In Fig. 11, we generate an audio clip of a talking man and another clip of police activities. The cascaded composition yields chaotic sounds as the police event context is not perceived. Our SSV2A composition successfully picks up this global cue and generates a police officer’s voice followed by a gun loading/shooting sound to indicate police events.
@@ A.6.3 Style Awareness
Figure 13: Screenshot of web demo application. Please visit our website and click the demo button to access this interface.
Appendix A.7 Future Improvements on Limitations
A.7.1 Temporal Synchronization
Our simple TA module has shown decent synchronization capability by attending to global audio scene semantics, as evidenced by WMAO evaluations in the Tab. 3. In reality, each sound source can have different temporal “activation” intervals in an audio scene. For example, a dog may bark in only the first three seconds of a video, followed by a baby laughing. In future works, we propose to perceive each sound source’s activations locally in time, similar to the tracklet detection [10] concept in video multi-object tracking. With this finer-grained temporal aggregation approach, we aim to further enhance SSV2A’s temporal alignment performance.
A.7.2 Sensitivity of Audio Conditioning
As stated in the limitations section, the lower sensitivity of our method’s audio conditioning compared to other modalities’ input is due to the lack of CLIP semantics for audio sound source inputs. An existing method, Wav2CLIP [63], translates CLAP audio embeddings to highly relevant CLIP embeddings. However, it operates on ViT-B/32 instead of ViT-L/14, which is the CLIP variant SSV2A employs. In future works, we plan to train a Wav2CLIP model compatible with SSV2A to address the audio sensitivity issue.
@@ Appendix A.8 Implementation
We upload the implementation and pretrained weights of SSV2A to github at https://github.com/wguo86/SSV2A. Please follow the readme instructions to set up a virtual environment and execute the designated scripts.
Appendix A.9 Web Demo Instructions
We implement a web application for interactive demonstration, as illustrated in Fig. 13. The left panel supports entries of generation conditions from image, text and audio. The right panel offers further hyperparameter tuning. We host this application on our own machine via a secure internet tunnel, strictly avoiding any collection of user data.